What are the mole fractions of hydrochloric acid (HCI) and water in a 20% (w/w) aqeous HCl solution?

1
Expert's answer
2022-03-13T15:20:52-0400

Solution:

Suppose we are given 100 g of an aqueous HCl solution

Therefore,

Mass of HCl = Mass of solution × w(HCl) / 100% = (100 g × 20%) / (100%) = 20 g

Mass of H2O = Mass of solution − Mass of HCl = 100 g − 20 g = 80 g


The molar mass of HCl is 36.458 g/mol

The molar mass of H2O is 18.015 g/mol


Calculate the moles of each component:

Moles of HCl = (20 g HCl) × (1 mol HCl / 36.458 g HCl) = 0.5486 mol HCl

Moles of H2O = (80 g H2O) × (1mol H2O / 18.015 g H2O) = 4.4407 mol H2O

Therefore,

Total moles of solution = Moles of HCl + Moles of H2O = 0.5486 mol + 4.4407 mol = 4.9893 mol


Calculate the mole fraction (χ) of each component:

χ = Moles of component / Total moles of solution

Therefore,

χHCl = Moles of HCl / Total moles of solution = (0.5486 mol) / (4.9893 mol) = 0.1099 = 0.11

χHCl = 0.11

χH2O = Moles of H2O / Total moles of solution = (4.4407 mol) / (4.9893 mol) = 0.8900 = 0.89

χH2O = 0.89


Answer:

The mole fraction of hydrochloric acid (HCI) is 0.11

The mole fraction of water (H2O) is 0.89